# Towards Imperceptible JPEG Deep Hiding: Multi-range Representations-driven Adversarial Stego Generation

### v1

@@ Abstract
Deep hiding has been exploring the hiding capability of deep learning-based models, aiming to conceal image-level messages into cover images and reveal them from generated stego images. Existing schemes are easily detected by steganalyzers due to their large payloads and their limitation to feature extraction based solely on either pure convolution or pure transformer operators within a single range, as well as pixel-level loss constraints. To address the issue, in this paper, we introduce generation-based adversarial attacks into color JPEG image deep hiding and propose a multi-range representations-driven adversarial stego generation framework called MRAG from a steganalysis perspective. Specifically, we integrate the local-range neighbor reception characteristic of the convolution and the global-range dependency modeling of the transformer to construct MRAG. Meanwhile, we use the transformed images obtained through coarse-grained and fine-grained frequency decomposition as inputs, introducing multi-grained information. Furthermore, a features angle-norm disentanglement loss is designed to constrain the generated stegos closer to covers in the angle and norm space of the steganalyzer’s classified features. Consequently, small yet effective adversarial perturbations can be injected into the process of generating stegos, ensuring that stegos maintain favorable secret restorability and imperceptibility. Extensive experiments demonstrate that MRAG can achieve state-of-the-art performance.
1 Introduction
@@
Image steganography is the art and science of sheltering confidential information in publicly available cover images to obtain stego ones, from which secret information can be revealed, and resulting stegos should be as unsuspicious as possible Provos and Honeyman (2003). It usually has three primary metrics Pevný and Fridrich (2008): the capacity refers to embedded payloads, the visual imperceptibility regarding the cover and stego being indistinguishable by naked eyes, and the steganalysis imperceptibility against the steganalyzers. Traditional steganographic methods and earlier deep learning-based methods maintain the bit-level payloads, typically setting it below 0.4 bits per pixel (BPP), with traditional methods going up to 1 BPP at most, and deep learning-based methods reaching up to 4.5 BPP Lu et al. (2021a). In light of this, Baluja Baluja (2017) delves deeper into the potential of steganographic models based on deep learning, successfully hiding a spatial RGB image into another one of the same size, known as deep hiding. Three natural advantages of deep hiding are also pointed out: 1) Since the type of hidden secret information is an image with high inherent correlations, the difficulty of network learning can be appropriately reduced. 2) The secret itself has a certain degree of fault tolerance. 3) The magnitude of the payload can be up to 24 BPP with a 1:1 ratio between the hidden secret image and the cover image. The success of the pioneering work, along with three natural advantages, paves the way for the flourishing of deep hiding.
The act of hiding inevitably yields embedded artifacts within the stegos. Whether artifacts are discovered usually depends on three factors Baluja (2020): the hidden capacity, the imperceptibility of the algorithm, and the texture of the image itself. The likelihood of the hiding behavior being discovered is negatively correlated with the latter two factors, and positively correlated with the capacity. In this sense, the capacity advantage of deep hiding also makes it more susceptible to detection, turning it into a disadvantage. Deep hiding with spatial RGB images as covers, abbreviated as DH-Spatial, has made significant progress recently Zhang et al. (2020); Lu et al. (2021a); Xu et al. (2022). However, the latest work, EFDR Yang and Liao (2023), points out that DH-Spatial methods perform poorly when the cover is the ubiquitous lossy image format - color JPEG image, resulting in visible color distortion, blocking, blurring, and banding artifacts in the generated JPEG stego. Similar visible artifacts also appear in the restored secret revealed from the JPEG stego, as shown in Fig.1 (a). Although EFDR achieves acceptable visual imperceptibility and restorability, its steganalysis imperceptibility remains unsatisfactory, as depicted in Fig.1 (b). The common shortcomings of these schemes are that they rely solely on single-range feature extraction using either pure convolution or pure transformer operators, and only consider pixel-level loss constraints.
A steganalyzer is a binary classifier that distinguishes between cover and stego images. With the development of deep learning, the way of features extraction has changed from manual to automatic extraction by networks and has become mainstream Wei et al. (2024); Hu and Wang (2024). Deep learning-based classifiers are found to be vulnerable to adversarial examples that can cause misclassification, which brings the prospect of resisting steganalyzers based on deep learning Zhang et al. (2018b). Reference Szegedy et al. (2014) puts forward that the adversarial example can be viewed as “accidental steganography”. They are achieved by adding low amplitude signals to an image for adversarial attacks or secret embedding and extraction. Analogously, deep hiding and generation-based adversarial attacks employ an image-to-image generative architecture to learn the image transformation task from the input image to the generated image Liu and Hsieh (2019). Therefore, in this paper, we attempt to introduce generation-based adversarial attacks into color JPEG image deep hiding, exploring the potential of generated color JPEG stego images in terms of secret restorability and imperceptibility, as illustrated in Fig.1 (c).
@@
For the steganalyzer, to better capture subtle changes, it is essential to consider both local-range and global-range information Wei et al. (2024). This in turn inspires us to integrate the local-range neighbor reception characteristic of convolution and the global-range dependency modeling of transformer, to construct a multi-range representations-driven adversarial stego generation framework called MRAG from a steganalysis perspective. Specifically, MRAG comprises four key components: the invertible local-range branch, the invertible global-range branch, the adaptive fusion module, and the surrogate steganalyzer. We also design the coarse-grained frequency decomposition (CFD) and the fine-grained frequency decomposition (FFD) to transform the input. According to the dimension characteristic of the obtained transformed images, the transformed images are used as the input for the local and global branches, respectively, naturally introducing multi-grained frequency information. Moreover, a features angle-norm disentanglement loss is designed to disentangle the surrogate steganalyzer’s classified features into the angle space associated with inter-class semantic differences and the norm space related to intra-class confidence, respectively. It can constrain the generated stegos closer to covers by pulling together the angle between the classified features of cover and stego, while keeping the norm changeless. With the help of features angle-norm disentanglement loss, small yet effective adversarial perturbations can be injected based on the multi-range representations extracted by local and global branches, ensuring that stegos maintain favorable secret restorability and imperceptibility. Our contributions can be summarized as follows:
We propose a multi-range representations-driven adversarial stego generation framework, MRAG, to generate color JPEG stego images with favorable secret restorability and imperceptibility. To our knowledge, this is the first to focus on the steganalysis imperceptibility issue of color JPEG image deep hiding.
We integrate the local-range neighbor reception of the convolution and the global-range dependency modeling of the transformer to extract multi-range representations from a steganalysis perspective, injecting adversarial perturbations based on these representations.
A features angle-norm disentanglement loss is designed to constrain the angle and norm distances between the steganalyzer’s classified features of the cover and the stego, in which the stego can be closer to the cover.
@@
Experiments validate the effectiveness of MRAG, demonstrating that MRAG establishes a new state-of-the-art benchmark in color JPEG image deep hiding.
2 Related Work
In this section, we summarize advancements in deep hiding and briefly introduce generation-based adversarial attacks.
2.1 Deep hiding
@@
Seminal work Baluja (2017) designs a deep hiding model that contains a secret pre-processing module, a hiding module, and a revealing module. UDH Zhang et al. (2020) is a cover-agnostic deep hiding model, which discards the pre-processing structure and only encodes secret images in the hiding sub-network. ISN Lu et al. (2021a) treats embedding and extraction as an inverse problem and utilizes the reversibility of invertible neural networks to address this question through its forward and backward processes. HiNet Jing et al. (2021) combines invertible neural networks with the wavelet domain and introduces low-frequency wavelet loss. RIIS Xu et al. (2022), a robust deep hiding model, takes various distortions into consideration, which shows acceptable robustness. The cover types and operators involved in the aforementioned works are always spatial RGB images and pure convolution. EFDR Yang and Liao (2023) is the first to research hiding image-level information in color JPEG images, focusing on mining fine-grained DCT representations using the pure transformer operator. Although advances in deep hiding have been remarkable, existing schemes are limited to single-range feature extraction, relying solely on either pure convolution or pure transformer operators. Moreover, these schemes are constrained only by pixel-level loss, leading to poor imperceptibility. It should be noted that, in this paper, we also perform deep hiding in color JPEG images, the ubiquitous lossy format images in daily life, and pay attention to its steganalysis imperceptibility issue first.
2.2 Generation-based adversarial attacks
Reference Szegedy et al. (2014) unveils that deep learning-based classifiers are susceptible to adversarial examples. Adding tiny perturbations to the input can cause misclassification with high confidence. Adversarial attacks can mainly be divided into two categories: gradient-based attacks and generation-based attacks Macas et al. (2024). Generation-based attacks usually utilize an image-to-image generative architecture to learn the mapping from input images to adversarial ones. Once the adversarial generator is trained well, only one forward propagation is needed to produce plenty of adversarial examples quickly. For example, AdvGAN Xiao et al. (2018), a representative of generation-based attacks, consists of an adversarial generator, a discriminator that differentiates between the original and the adversarial example, and a surrogate classifier intended to be attacked. However, SSAE Lu et al. (2021b) demonstrates that the discriminator is redundant if perturbations are tiny enough. Therefore, our MRAG also discards the discriminator to simplify the model.
3 Proposed Framework
3.1 Overview
@@
The pipeline of MRAG consists of four key components: the invertible local-range branch, the invertible global-range branch, the adaptive fusion module, and the surrogate steganalyzer, as shown in Fig. 2. Since we construct MRAG utilizing invertible neural networks as much as possible to ensure reversibility, it can implement forward-hiding and backward-revealing processes with shared parameters. The other image conversion operations and network structures are also reversible and symmetric. Conversion operations contain the coarse-grained frequency decomposition (CFD), the fine-grained frequency decomposition (FFD), and the inverse FFD (IFFD). The input of MRAG’s forward-hiding process is the color JPEG cover image and the spatial RGB secret image . The input is converted into coarse-grained and fine-grained frequency maps through CFD and CFDFFD, respectively. According to the dimension characteristic of the obtained frequency maps, they are fed into the local and global branches’ forward processes, respectively. Then, the global branch’s output passes through the IFFD and is concatenated with the local branch’s output. Afterward, their concatenation is fed into the adaptive fusion module’s forward process to obtain the JPEG stego image and the redundant information , an all-zero matrix. MRAG’s backward-revealing process takes the color JPEG stego image and as input to reveal the secret image. For effectively inducing the surrogate steganalyzer to misclassify the cover and stego , we conduct the features angle-norm disentanglement and pull together the angles between the classified features of cover and the classified features of stego while keeping their norm values changeless.
3.2 MRAG Model
@@
Given that invertible neural networks can preserve as much detail as possible due to the theoretically lossless information characteristic Dinh et al. (2014, 2017); Gilbert et al. (2017), we utilize invertible neural networks to construct our MRAG to the greatest extent. As shown in Fig. 2, several key components of MRAG, including the invertible local-range branch, the invertible global-range branch, and the adaptive fusion module, all employ invertible neural networks. For the basic unit of invertible neural networks, the affine coupling layer, the input needs to be divided into two parts, and , along the channel axis. Following reference Xiao et al. (2020), we use an additive affine transformation for the upper part, , and employ an enhanced affine transformation for the lower part, . To be specific, the forward process of the affine coupling layer can be formulated as
@@
Convolution performs excellently in image-related tasks and has nice local neighbor reception characteristics Krizhevsky et al. (2012). Here, , , and in the invertible local-range branch is implemented by a 5-layer dense block with a convolution kernel Huang et al. (2017). In the dense block, the output feature maps of each layer are passed to all subsequent layers, greatly enhancing the ability for representation sharing and learning. For the invertible local-range branch, the input is coarse-grained frequency maps through CFD. CFD includes the Huffman decoding of the color JPEG cover image and the DCT transformation of the spatial RGB secret image.
@@
To further enhance representation and global dependency modeling ability, in the invertible global-range branch, , , and adopt a standard 1-layer Pre-LN transformer Dosovitskiy et al. (2021). This Pre-LN transformer applies the layer normalization before the multi-head self-attention and the multi-layer perceptron for a more stable gradient distribution. For the invertible global-range branch, the input is the fine-grained frequency maps obtained through CFDFFD. FFD includes the sub-band gathering and the zigzag scanning of the coarse-grained frequency maps. The sub-band gathering procedure rearranges coarse-grained frequency maps to place the frequency components of the same sub-band on the same channel, making frequency components more compact. The zigzag scanning procedure further rearranges them along the channel axis to group similar frequency sub-bands together, exploiting the structural information in the JPEG transformation.
@@
Our designed adaptive fusion module consists of a pair of convolutions with channel attention and a series of affine coupling layers with an invertible convolution. , , and in the adaptive fusion module also employ a 5-layer dense block. The convolution with channel attention designed in reference Tan et al. (2022) is used to adaptively extract the most informative frequency representations. The invertible convolution is a kind of convolution commonly used for invertible neural networks. Adding an invertible convolution before each affine coupling layer can increase the dependencies between representations, allowing for more effective information fusion. For the adaptive fusion module, the input is the concatenation of representations extracted from the invertible local-range branch and representations extracted from the invertible global-range branch after undergoing IFFD. IFFD is used for dimension alignment.
@@
In our MRAG, the surrogate steganalyzer can be any well-trained JPEG steganalyzer. Here, we select three open-source steganalyzers: UCNet Wei et al. (2022), EWNet Su et al. (2021), and DBS2Net Hu and Wang (2024). UCNet is the sole steganalyzer tailored for color JPEG images. The others are straightforwardly based on existing steganalyzers for gray-scale JPEG images or on some well-known models in computer vision. EWNet and DBS2Net are gray-scale JPEG steganalyzers. Here, following the setting in reference Wei et al. (2022), we adjust the number of input channels in the first convolution layer to 3, thereby enabling them to detect color images.
@@
ISN+ Lu et al. (2021a) | 19.85 | .662 | 17.52 | .344 | 19.81 | .643 | 18.98 | .361 | 22.55 | .731 | 14.60 | .345 |
RIIS Xu et al. (2022) | 24.98 | .798 | 9.28 | .169 | 24.97 | .781 | 9.80 | .178 | 29.20 | .864 | 5.84 | .143 |
EFDR Yang and Liao (2023) | 33.95 | .918 | 4.31 | .028 | 34.05 | .919 | 4.37 | .027 | 36.43 | .921 | 3.22 | .032 |
MRAG (UCNet) | 40.11 | .983 | 1.88 | .004 | 39.85 | .980 | 2.04 | .004 | 41.54 | .975 | 1.66 | .005 |
MRAG (EWNet) | 41.67 | .988 | 1.62 | .001 | 41.20 | .985 | 1.80 | .001 | 42.32 | .978 | 1.55 | .002 |
MRAG (DBS2Net) | 37.14 | .978 | 2.78 | .006 | 37.00 | .973 | 2.90 | .005 | 39.49 | .969 | 2.15 | .007 |
@@
ISN+ Lu et al. (2021a) | 24.09 | .828 | 9.96 | .208 | 24.08 | .823 | 10.66 | .213 | 27.12 | .830 | 7.80 | .239 |
RIIS Xu et al. (2022) | 28.25 | .897 | 5.63 | .098 | 28.56 | .895 | 5.86 | .104 | 32.76 | .921 | 3.93 | .102 |
EFDR Yang and Liao (2023) | 32.22 | .927 | 4.71 | .025 | 32.04 | .924 | 4.93 | .027 | 36.01 | .944 | 3.12 | .027 |
MRAG (UCNet) | 31.68 | .920 | 5.01 | .028 | 31.53 | .916 | 5.24 | .031 | 35.60 | .940 | 3.27 | .031 |
MRAG (EWNet) | 32.78 | .936 | 4.39 | .023 | 32.62 | .934 | 4.58 | .025 | 36.37 | .950 | 2.98 | .026 |
MRAG (DBS2Net) | 32.35 | .930 | 4.65 | .022 | 32.16 | .927 | 4.86 | .024 | 35.92 | .945 | 3.16 | .026 |
3.3 Loss Function
@@
Several works Liu et al. (2018); Chen et al. (2020) conjecture that classified features learned by classifiers can naturally disentangle into angle and norm distances related to inter-class semantic differences and intra-class confidence, respectively. The steganalyzer can be viewed as a binary classifier. The classified features learned by the steganalyzer are crucial for distinguishing between covers and stegos. Inspired by these facts, we introduce this loss to constrain the generated stegos closer to covers in the angle and norm space of the steganalyzer’s classified features, thereby enhancing the steganalysis imperceptibility. For fooling the surrogate steganalyzer, we focus on pulling together the angles between the classified features of cover and the classified features of stego . Additionally, to minimize the injected adversarial perturbations, we keep the norm values of and changeless. Thus, we can define the features angle-norm disentanglement loss as follows
where indicates the cosine similarity. measures the similarity between two feature vectors by the cosine value of the angle between them. A larger value indicates a smaller angle, and hence a higher similarity. is the mean squared error. and denote the angle loss and norm loss of the classified features, respectively.

### v2

# Towards Imperceptible JPEG Image Hiding: Multi-Range Representations-Driven Adversarial Stego Generation

@@ Abstract
Image hiding fully explores the hidden potential of deep learning-based models, aiming to conceal image-level messages within cover images and reveal them from stego images to achieve covert communication. Existing hiding schemes are easily detected by the naked eyes or steganalyzers due to the cover type confined to the spatial domain, single-range feature extraction and attacks, and insufficient loss constraints. To address these issues, we propose a multi-range representations-driven adversarial stego generation framework called MRAG for JPEG image hiding. This design stems from the fact that steganalyzers typically combine local-range and global-range information to better capture hidden traces. Specifically, MRAG integrates the local-range characteristic of the convolution and the global-range modeling of the transformer. Meanwhile, a features angle-norm disentanglement loss is designed to launch multi-range representations-driven feature-level adversarial attacks. It computes the adversarial loss between covers and stegos based on the surrogate steganalyzer’s classified features, i.e., the features before the last fully connected layer. Under the dual constraints of features angle and norm, MRAG can delicately encode the concatenation of cover and secret into subtle adversarial perturbations from local and global ranges relevant to steganalysis. Therefore, the resulting stego can achieve visual and steganalysis imperceptibility. Moreover, coarse-grained and fine-grained frequency decomposition operations are devised to transform the input, introducing multi-grained information. Extensive experiments demonstrate that MRAG can achieve state-of-the-art performance.
Introduction
@@
Image steganography is the art and science of sheltering confidential information in publicly available cover images to obtain stego ones. The resulting stegos should be as unsuspicious as possible, from which secret can be revealed (Provos and Honeyman 2003). It typically has three primary metrics (Pevný and Fridrich 2008): the capacity regarding embedded payloads, the visual imperceptibility by the naked eyes, and the steganalysis imperceptibility against the steganalyzers. The steganalyzer is a binary classifier that distinguishes stego images from cover images. Traditional steganographic methods and earlier deep learning-based methods maintain the bit-level payloads, typically below 0.4 bits per pixel (BPP). The traditional can reach 1 BPP at most, while the deep learning-based can go as high as 4.5 BPP (Lu et al. 2021). In light of this, Baluja (Baluja 2017) explores the potential of deep learning-based steganographic models, successfully embedding a spatial RGB image into another of the same size, known as image hiding. Three natural advantages of image hiding are also pointed out: 1) Since the secret is an image with high inherent correlations, the difficulty of network learning can be reduced. 2) The secret itself has a certain fault tolerance. 3) The payload can reach 24 BPP with a 1:1 ratio between the secret and cover. The success of this pioneering work, along with three natural advantages, facilitates the flourishing development of image hiding.
The act of hiding inevitably yields embedded artifacts. Whether artifacts are discovered usually depends on three factors (Baluja 2020): the hidden capacity, the imperceptibility of the algorithm, and the texture of the image itself. The likelihood of detection is negatively correlated with the latter two factors, but positively correlated with the hidden capacity. In this sense, the capacity advantage of image hiding also makes it more vulnerable to detection, turning into a disadvantage. Most hiding methods (Zhang et al. 2020; Lu et al. 2021; Jing et al. 2021; Guan et al. 2023; Xu et al. 2022) default to such a disadvantage, preferring to improve capacity, visual imperceptibility, and robustness. Only several works (Zheng et al. 2023; Zhang et al. 2021; Hu et al. 2024) explore the steganalysis imperceptibility utilizing generative adversarial networks (GAN) or adversarial attacks. Due to the inherent instability and convergence difficulties in min-max optimization, GAN-based approaches perform poorer steganalysis imperceptibility. To further improve steganalysis imperceptibility, adversarial attacks-based methods typically employ a pre-trained surrogate steganalyzer and introduce an adversarial loss. Nevertheless, the cover type of the aforementioned methods is exclusively lossless spatial RGB images. We term these methods as spatial image hiding (SIH) for brevity. When the cover is a color JPEG image, the ubiquitous lossy image format, SIH methods perform poorly. There are visible artifacts in the generated JPEG stego and the restored secret. Even the robust SIH method that resists JPEG compression, RIIS (Xu et al. 2022), also encounters the same problem, as shown in Fig.1 (a).
Additionally, the existing SIH methods are restricted to single-domain feature extraction and attacks, lacking steganalysis domain knowledge. The adversarial loss is typically computed as cross-entropy loss based on the surrogate steganalyzer’s output. The recent EFDR (Yang and Liao 2023) is currently the only work that uses color JPEG images as the covers. Although EFDR can effectively mitigate visible artifacts, its steganalysis imperceptibility remains unsatisfactory due to single-range feature extraction and insufficient loss constraints, as depicted in Fig.1 (b). In summary, existing hiding schemes are easily detected by the naked eyes or steganalyzers due to the cover type confined to the spatial domain, single-range feature extraction and attacks, and insufficient loss constraints. To address these issues, in this paper, we attempt to develop a novel adversarial stego generation framework for color JPEG image hiding, exploring the potential of the generated color JPEG stego in terms of secret restorability, visual imperceptibility, and steganalysis imperceptibility, as illustrated in Fig.1 (c).
@@
For the steganalyzer, to better capture subtle changes, it is essential to consider both local-range and global-range information (Wei, Luo, and Huang 2024). This, in turn, inspires us to construct a multi-range representations-driven adversarial stego generation framework called MRAG from a steganalysis perspective. MRAG integrates the local-range characteristic of the convolution and the global-range modeling of the transformer. Specifically, MRAG comprises four key components: the invertible local-range branch, the invertible global-range branch, the adaptive fusion module, and the surrogate steganalyzer. A features angle-norm disentanglement loss is incorporated to launch multi-range representations-driven feature-level adversarial attacks. It computes the adversarial loss between covers and generated stegos based on the surrogate steganalyzer’s classified features. Therefore, MRAG is constrained to delicately encode the concatenation of cover and secret into subtle adversarial perturbations from local and global perspectives relevant to steganalysis. We also tailor the coarse-grained frequency decomposition (CFD) and the fine-grained frequency decomposition (FFD) to transform the input. According to the dimension of the obtained transformed images, the transformed images are used as the input for the local and global branches, respectively, naturally introducing multi-grained frequency information. Our contributions can be summarized as follows:
We propose a novel adversarial stego generation framework called MRAG for color JPEG image hiding. MRAG integrates the local-range characteristic of the convolution and the global-range modeling of the transformer from a steganalysis perspective. To our best knowledge, this is the first attempt to explore the potential of the generated color JPEG stego in terms of secret restorability, visual imperceptibility, and steganalysis imperceptibility.
We design a features angle-norm disentanglement loss based on the surrogate steganalyzer’s classified features, which is incorporated to launch multi-range representations-driven feature-level adversarial attacks. Under the dual constraints of features angle and norm, MRAG can delicately encode the concatenation of cover and secret into subtle adversarial perturbations from local and global ranges relevant to steganalysis.
@@
Comprehensive experiments validate the effectiveness and superiority of MRAG on three widely-used datasets, demonstrating that MRAG establishes a new state-of-the-art benchmark in color JPEG image hiding.
Related Works
In this section, we summarize advancements in image hiding and introduce features angle-norm disentanglement.
Image Hiding
@@
Seminal work (Baluja 2017) designs an image hiding model containing a secret pre-processing module, a hiding module, and a revealing module. UDH (Zhang et al. 2020) is a cover-agnostic image hiding model that discards the pre-processing structure and solely encodes secret images. ISN (Lu et al. 2021) treats embedding and extraction as an inverse problem and utilizes the forward and backward processes of invertible neural networks to address it. To improve visual imperceptibility, HiNet (Jing et al. 2021) combines invertible neural networks with the wavelet domain and introduces low-frequency wavelet loss. DeepMIH (Guan et al. 2023) proposes a novel framework for multiple image hiding. It designs an importance map module to guide the concealing process of the next secret image. CAIS (Zheng et al. 2023) devises a composition estimation network as the discriminator and promotes steganalysis imperceptibility through generative adversarial training. Both USAP (Zhang et al. 2021) and DIH-OAIN (Hu et al. 2024) utilize adversarial attacks to boost steganalysis imperceptibility. RIIS (Xu et al. 2022) is a robust image hiding method that considers various distortions. Nevertheless, the aforementioned works have common shortcomings. The cover type is always lossless spatial RGB images. Their model design lacks steganalysis domain knowledge, relying solely on single-domain feature extraction and attacks. Additionally, the adversarial loss is typically computed as cross-entropy loss based on the surrogate steganalyzer’s output. EFDR (Yang and Liao 2023) is the first to research image hiding in color JPEG images, focusing on mining fine-grained DCT representations using the transformer operator. Although EFDR has favorable visual imperceptibility and secret restorability, its steganalysis imperceptibility remains poor due to single-range feature extraction and insufficient loss constraints. Noted that, in this paper, we also perform image hiding in color JPEG images, the ubiquitous lossy image format. This is the first attempt to explore the potential of the generated color JPEG stego in terms of secret restorability, visual imperceptibility, and steganalysis imperceptibility.
Features Angle-Norm Disentanglement
References (Liu et al. 2018; Chen et al. 2020) unveil that the features before the last fully connected layer of the classifier, abbreviated as classified features, can naturally disentangle into angle and norm distances related to inter-class semantic differences and intra-class confidence, respectively. Specifically, they train a plain CNN classifier on the MNIST dataset. Then set the output dimension of classified features to 2 and visualize these features. Three key observations emerge from the visualization. 1) A well-trained classifier can effectively cluster features belonging to the same category. 2) Features angle accounts for semantic/label differences. 3) Features norm reflects intra-class variations. Inspired by these facts, we introduce features angle-norm disentanglement to compute adversarial loss, which pulls together the angles between the classified features of cover and stego while keeping their norm values changeless.
Proposed Framework
Overview
The pipeline of MRAG consists of four key components: the invertible local-range branch, the invertible global-range branch, the adaptive fusion module, and the surrogate steganalyzer, as shown in Fig. 2. We construct MRAG utilizing invertible neural networks, the mainstream learning paradigm for image hiding, to ensure reversibility as much as possible. The other image conversion operations and network structures are also reversible and symmetric. Therefore, we can implement hiding and revealing utilizing MRAG’s forward and backward processes.
@@
The inputs of MRAG’s forward-hiding process are the color JPEG cover image and the spatial RGB secret image . Conversion operations contain the coarse-grained frequency decomposition (CFD), the fine-grained frequency decomposition (FFD), and the inverse FFD (IFFD). The inputs are converted into coarse-grained and fine-grained frequency maps through CFD and CFDFFD, respectively. According to the dimension characteristic of the obtained frequency maps, they are fed into the local and global branches’ forward processes, respectively. The dimensions of coarse-grained and fine-grained frequency maps are and respectively. The superscript denotes the shape with . The dimension of fine-grained frequency maps is better suited for the patch-wise architecture of the transformer. Then, the global branch’s output passes through the IFFD and is concatenated with the local branch’s output. Afterward, the concatenation is fed into the adaptive fusion module’s forward process to obtain the JPEG stego image and the redundant information , an all-zero matrix. MRAG’s backward-revealing process takes the color JPEG stego image and as input to reveal the secret image. For effectively inducing the surrogate steganalyzer to misclassify the cover and stego , we conduct the features angle-norm disentanglement and pull together the angles between the classified features of cover and the classified features of stego while keeping their norm values changeless.
MRAG Model
@@
Given that invertible neural networks, the mainstream learning paradigm for image hiding, can preserve as much detail as possible due to their theoretical information lossless characteristic (Dinh, Krueger, and Bengio 2014; Dinh, Sohl-Dickstein, and Bengio 2017; Gilbert et al. 2017), we utilize them to construct our MRAG to the greatest extent. As shown in Fig. 2, key components of MRAG, including the invertible local-range branch, the invertible global-range branch, and the adaptive fusion module, employ invertible neural networks. For the affine coupling layer, the basic unit of invertible neural networks, the input needs to be divided into two parts, and , along the channel axis. Following the reference (Xiao et al. 2020), we use an additive affine transformation for the upper part, , and employ an enhanced affine transformation for the lower part, . To be specific, the forward process of the affine coupling layer can be formulated as
@@
As we know, convolution performs excellently in image-related tasks and has nice local neighbor reception characteristics (Krizhevsky, Sutskever, and Hinton 2012). Here, in the invertible local-range branch, , , and are implemented by a classical 5-layer dense block with a convolution kernel (Huang et al. 2017). In the dense block, the output feature maps of each layer are passed to subsequent layers, greatly enhancing representation sharing and learning. For the invertible local-range branch, the input is coarse-grained frequency maps through CFD. CFD includes the Huffman decoding of the color JPEG cover image and the DCT transformation of the spatial RGB secret image.
@@
To further enhance representation and global dependency modeling, , , and in the invertible global-range branch adopt a standard 1-layer Pre-LN transformer (Dosovitskiy et al. 2021). The Pre-LN transformer applies the layer normalization before the multi-head self-attention and the multi-layer perceptron for a more stable gradient distribution. For the invertible global-range branch, the input is the fine-grained frequency maps obtained through CFDFFD. FFD includes the sub-band gathering and the zigzag scanning. The sub-band gathering rearranges coarse-grained frequency maps, placing the frequency components of the same sub-band on the same channel, making frequency components more compact. The zigzag scanning further rearranges them along the channel axis, grouping similar frequency sub-bands and exploiting the structural information in the JPEG transformation.
@@
In addition, to enhance the interaction between local and global representations extracted by the aforementioned two branches, the proposed adaptive fusion module consists of a pair of convolutions with channel attention and a series of affine coupling layers with an invertible convolution. , , and in the adaptive fusion module also employ the classical 5-layer dense block. The convolution with channel attention, designed in the reference (Tan et al. 2022), is used to extract the most informative frequency representations adaptively. Adding an invertible convolution before each affine coupling layer can increase the dependencies between representations, enabling more effective information fusion. For the adaptive fusion module, the input is the concatenation of representations extracted from the invertible local-range branch and representations extracted from the invertible global-range branch after undergoing IFFD. IFFD is used for dimension alignment.
@@
The surrogate steganalyzer in MRAG can be any well-trained JPEG steganalyzer. Here, we select three open-source steganalyzers: UCNet (Wei et al. 2022), EWNet (Su, Zhao, and He 2021), and DBS2Net (Hu and Wang 2024). UCNet is the sole steganalyzer tailored for color JPEG images. The others are straightforwardly based on existing steganalyzers for gray-scale JPEG images or some well-known computer vision models. EWNet and DBS2Net are gray-scale JPEG steganalyzers. Here, following the setting in the reference (Wei et al. 2022), we adjust the number of input channels in the first convolution layer from 1 to 3, enabling them to detect color images.
@@
| ISN+ (Lu et al. 2021) | 19.85 | .662 | 17.52 | .344 | 19.81 | .643 | 18.98 | .361 | 22.55 | .731 | 14.60 | .345 |
| RIIS (Xu et al. 2022) | 24.98 | .798 | 9.28 | .169 | 24.97 | .781 | 9.80 | .178 | 29.20 | .864 | 5.84 | .143 |
| EFDR (Yang and Liao 2023) | 33.95 | .918 | 4.31 | .028 | 34.05 | .919 | 4.37 | .027 | 36.43 | .921 | 3.22 | .032 |
[…]
@@
| ISN+ (Lu et al. 2021) | 24.09 | .828 | 9.96 | .208 | 24.08 | .823 | 10.66 | .213 | 27.12 | .830 | 7.80 | .239 |
| RIIS (Xu et al. 2022) | 28.25 | .897 | 5.63 | .098 | 28.56 | .895 | 5.86 | .104 | 32.76 | .921 | 3.93 | .102 |
| EFDR (Yang and Liao 2023) | 32.22 | .927 | 4.71 | .025 | 32.04 | .924 | 4.93 | .027 | 36.01 | .944 | 3.12 | .027 |
[…]
Loss Function
@@
For fooling the surrogate steganalyzer, we focus on pulling together the angles between the classified features of cover and the classified features of stego . Additionally, to minimize the adversarial perturbations, we keep the norm values of and changeless. Thus, we can define the features angle-norm disentanglement loss as follows
where indicates the cosine similarity. measures the similarity between two feature vectors by the cosine value of the angle between them. A larger value indicates a smaller angle, hence a higher similarity. is the mean squared error. and denote the angle loss and norm loss of the classified features, respectively.